OpenStudy (anonymous):

FAN AND MEDAL! 70 is what percent of 1000?

OpenStudy (acxbox22):

70/1000=x/100 7/100=x/100 x=?

OpenStudy (anonymous):

7%

OpenStudy (anonymous):

thanks ^_^

OpenStudy (acxbox22):

;)